Walkenhorst Company’s machining department prepared its 2019 budget based on the following data:

Practical capacity 40,000 units
Standard machine hours per unit 2
Standard variable factory overhead $3.00 per machine hour
Budgeted fixed factory overhead $ 296,000

The department uses machine hours to apply factory overhead to production. In 2019, the department used 86,300 machine hours and incurred $565,000 in total manufacturing overhead cost to manufacture 42,120 units. Actual fixed overhead cost for the year was $308,000.

Required:

Determine for the year:

1. The fixed, variable, and total factory overhead application rates (per machine hour). (Round your answers to 2 decimal places.)

2. The total flexible budget, for factory overhead cost based on output achieved in 2019.

3. The production volume variance. State whether this variance was favorable (F) or unfavorable (U).

4. The total overhead spending variance. State whether this variance was favorable (F) or unfavorable (U).

5. The overhead efficiency variance. State whether this variance was favorable (F) or unfavorable (U).

6. The variable overhead spending variance and the fixed overhead spending variance. State whether each variance is favorable (F) or unfavorable (U).
Answer #1

Solution 1:

Fixed overhead rate = Budgeted fixed overhead / Budgeted machine hours = $296,000 / 80000 = $3.70 per machine hour

Variable factory overhead rate = $3 per machine hour

Total factory overhead application rate = $3.70 + $3 = $6.70 per machine hour

Solution 2:

Budgeted variable overhead for actual output = 42120*2*$3 = $252,720

Budgeted fixed overhead fro actual output = $296,000

Total flexible budget for factory overhead cost = $252,720 + $296,000 = $548,720

Solution 3:

Production volume variance = Fixed overhead applied - Budgeted fixed overhead

= (42120*2*$3.70) - $296,000 = $15,688 F

Solution 4:

Total overhead spending variance = Total budgeted overhead - Total actual overhead

= $548,720 - $565,000 = $16,280 U

Note: I have answered first 4 parts of the question as HOMEWORKLIB RULES policy, kindly post separate question for answer of remaining parts.